# Spectrum from Defocus: Fast Spectral Imaging with Chromatic Focal Stack

### v1

@@ Abstract
Hyperspectral cameras face harsh trade-offs between spatial, spectral, and temporal resolution in an inherently low-photon regime. Computational imaging systems break through these trade-offs with compressive sensing, but require complex optics and/or extensive compute. We present Spectrum from Defocus (SfD), a chromatic focal sweep method that recovers state-of-the-art hyperspectral images with a small system of off-the-shelf optics and $<1$ second of compute. Our camera uses two lenses and a grayscale sensor to preserve nearly all incident light in a chromatically-aberrated focal stack. Our physics-based iterative algorithm efficiently demixes, deconvolves, and denoises the blurry grayscale focal stack into a sharp spectral image. The combination of photon efficiency, optical simplicity, and physical modeling makes SfD a promising solution for fast, compact, interpretable hyperspectral imaging.
Figure 1: Comparison and overview of our Spectrum from Defocus (SfD) camera. (a) Our camera provides state-of-the-art PSNR with simple optics and a high frame rate, which we compute as $1/(t_{\text{exposure}}+t_{\text{compute}})$ with a simulated 5 sec total exposure for each method. Bubble diameters correspond to the number of optical components in each method. See Table 1 for details. (b) Our hardware prototype uses a moving lens to change the focus of chromatically aberrated light on its way to a grayscale sensor. (c) Our imaging system captures a grayscale focal stack and reconstructs images in under a second. We show an RGB projection of a 12-channel recovered image, with a high-resolution inset (blue square) and reconstructed spectrum (red star).
@@ 1 Introduction
Hyperspectral imaging (HSI) extends traditional color imaging by capturing detailed spectral signatures across numerous spectral bands, rather than just the three broad RGB bands. This richer spectral information is essential for applications that rely on subtle material differences, such as remote sensing for environmental monitoring [3], medical diagnostics [20], food quality control [11], and material identification in industrial inspection [24]. Despite its importance, capturing high-quality spectral images with compact, affordable cameras remains challenging. Hyperspectral sensors suffer from inherently lower signal-to-noise ratios because incoming light is divided across many spectral channels, leaving each channel with reduced photon counts. This challenge is reflected in systems that rely on dispersive elements [32], which spread spectral information to the corresponding sensor pixels. This limitation make traditional hyperspectral cameras particularly fragile in low-light conditions or dynamic scenes that require short exposure times, limiting their practicality.
To mitigate this fragility, CASSI-like computational imaging systems have been introduced [28, 13, 19]. These systems encode spectral information into the spatial dimension using a coded aperture [28] and employ an optimization framework in the reconstruction phase. However, solving optimization problems in hyperspectral imaging entails high computational demands due to extensive matrix multiplications in iterative processes and may require large-scale matrix inversion. Although adopting a naive optimization algorithm (e.g., inverse filtering as used by Zhan et al.) can reduce computational complexity, it often degrades the quality of reconstructed images, limiting applicability to real-world data.
Data-driven methods [5, 21] have been proposed in conjunction with CASSI-like imaging systems for addressing the computational issues. These methods typically design parameterized modules to capture the intrinsic low-rank properties in both spatial and spectral dimensions. For instance, MST [5] introduces Spectral-wise Multi-head Self-Attention (S-MSA) to exploit spectral-wise similarity in hyperspectral image reconstruction. However, the hallucination [12] issue in data-driven models can restrict their applicability in tasks requiring precise spectral accuracy, such as material analysis and food quality control.
Considering these limitations—light inefficiency, high computational complexity, and hallucination—we propose a chromatic focal sweep camera consisting of two lenses and a grayscale sensor. The special sensing matrix structure in our physics-based iterative algorithm enables the reconstruction of high-quality and robust hyperspectral images. Furthermore, the physics-based approach mitigates the hallucination issue commonly observed in data-driven methods.
@@
  • Fast computation with a physics-based, interpretable model.
  • A simple optical design consisting of few, off-the-shelf (OTS) components, which outperforms systems with up to 20 optical elements.
Figure 2: Optical design. (a) The system consists of a lens pair in which the second lens is moved to $N$ positions, $z_{1},z_{2},\cdots,z_{N}$, respectively when capturing $N$ images of a scene. We assume the imaging target is within the depth of field throughout this paper. The $N$ captured images, $I_{1},\cdots,I_{N}$, individually focus $N$ discrete, predetermined wavelengths, $\lambda_{1},\cdots,\lambda_{N}$, while blurring others, forming a chromatic focal stack. (b) Measured point spread functions (PSFs) of the real-world prototype at different wavelengths $\lambda_{i}$ and lens position $z_{i}$. The PSFs clearly show the chromatic focal shift , as the focus shifts to different wavelengths at different lens positions $z_{i}$. Although six different lens positions are shown, only a subset of lens positions is used during reconstruction, corresponding to the number and location of captured measurements.
@@ 2 Related Works
Scanning Methods acquire hyperspectral data cubes by sequentially capturing individual 2D slices of the scene. The most straightforward implementation uses band-pass filters or liquid crystal tunable filters to capture one narrow spectral band per frame, gradually scanning across the spectrum [27]. Pushbroom systems take a slightly different approach, capturing a single spatial row at a time through a narrow slit, while dispersive optics spread each row into a full spectrum across the sensor [25]. While these techniques can deliver high spatial and spectral resolution, they suffer from low photon efficiency—since only a fraction of the incoming light is used for each measurement—and long acquisition times, which restricts them to static, well-lit scenes. In contrast, our method passively encodes spectral information into each frame and avoids light-blocking (low light transmission) elements entirely, enabling faster capture with higher photon efficiency.
@@
Compressive hyperspectral imaging systems capture optically coded measurements that must be computationally reconstructed to recover the final hyperspectral data cube. Among these, Coded Aperture Snapshot Spectral Imaging (CASSI) [28] was developed to overcome the limitations of sequential spectral scanning systems. A coded aperture is placed either before or after a dispersive prism in the optical path, encoding spectral signatures into the spatial domain. CASSI systems are typically divided into two categories based on their encoding strategy: spatially-encoded CASSI, which employs a single disperser (SD-CASSI) [28], and spatial-spectral CASSI, which codes information in both spatial and spectral domains (DD-CASSI [13], SS-CASSI [19]). Regardless of the encoding scheme, in all CASSI systems reconstruction algorithms play a critical role in determining final image quality. Early reconstruction methods relied on optimization-based solvers with hand-crafted priors [28], while more recent approaches introduced learning-based reconstructions [5, 22]. Hybrid methods further combined the system’s sensing matrix with deep learning, either to directly parameterize the solution space [9] or to learn data-driven priors for regularization [29]. Despite their success, traditional CASSI systems are often relatively large, and mechanically sensitive, making them impractical for compact imaging platforms. In contrast, our system requires only two off-the-shelf lenses and achieves spectral imaging through a focal sweep, significantly reducing hardware complexity.
@@
Another class of compressive spectral imagers relevant to this work are spectral-diffractive imagers, which encode spectral information using diffractive optical elements (DOEs) with optimized phase masks, followed by computational reconstruction to recover the hyperspectral data cube. Early work, such as [14], used simple diffraction gratings to simultaneously resolve spatial and spectral information, but these systems suffered from low resolution and bulky optical assemblies. Later, Heide et al. [16] introduced a co-designed pair of DOEs with optimized phase profiles, enabling tunable control over both focus and chromatic aberration without the need for large, conventional optics. They further proposed a blind deconvolution algorithm to jointly recover the scene and correct for encoding imperfections, improving robustness. More recently, efforts have focused on using a single DOE, reducing hardware complexity but introducing stronger optical artifacts which are typically handled through end-to-end deep learning reconstructions trained on large spectral datasets. Similar to these approaches, our system also relies on spectrally encoded PSFs; however, in contrast, we achieve this encoding using the naturally occurring chromatic aberration of standard refractive lenses, alleviating the need for custom-fabricated optics.
Recent Spectral Computational Imagers highlight a trade-off between optical simplicity and computational cost in spectral imaging. Spectral DiffuserCam (SDC) [23] uses a lensless design with a spectral filter array, but requires solving a computationally expensive inverse problem via FISTA. In contrast, KRISM [26] employs a complex optical system to directly measure the spectral singular values, offloading most of the computation to the hardware itself. Our work balances both approaches, combining a simple optical setup using off-the-shelf lenses with a lightweight, physics-based reconstruction.
Spectral Focal Sweeping and Stacking refers to techniques that capture images at multiple focus distances to jointly recover spectral, spatial, or depth information. [10] used uncorrected chromatic aberration to passively extend depth of field across wavelengths, allowing different colors to focus at different depths. [7] introduced a color-coded aperture, creating different depths of field for each RGB channel and using this as a cue for joint depth estimation and all-in-focus imaging. Both methods leverage chromatic defocus, but only for depth and RGB recovery. [15] modeled spatially-varying chromatic PSFs to correct color fringing in simple lenses, enhancing RGB image quality. In contrast, we preserve and directly exploit chromatic defocus to encode a full spectral cube across the visible range. Finally, [17] proposes a novel multispectral focal stack acquisition method. The system intentionally enhances chromatic aberration while suppressing other optical aberrations, allowing different spectral channels to focus on distinct scene depths enabling recovery of the full light field using an algorithm that fills in the missing details.
@@ 3 Method
Table 1: Comparison of SOTA hyperspectral imaging systems in terms of reconstruction performance, and computational and optical requirements. Reconstruction quality is benchmarked on 30 images from the Harvard dataset [6] and adopt a brightly lit condition with 5-second total exposure time and optical component effects (see supplement). Timings are reported for an NVIDIA RTX A6000. The count of optical components includes: lenses, apertures, prisms, actuators, SLMs, and sensors, but not control electronics.
Modality Reconstruction Performance Computation and Optics Category PSNR (dB) $\uparrow$ SSIM $\uparrow$ SAM () $\downarrow$ Compute Time FoV () Opt. Components TF -4.88 0.03 88.71 0 N.A. 3 Spectral Filter SDC [23] 12.03 0.37 40.88 $>$ 15 mins 14 3 MST [5] 21.18 0.53 38.87 1.19 s 10 9 Dispersive Element Choi et al. [9] 15.97 0.41 41.84 $>$ 15 mins 10 9 KRISM [26] 21.12 0.58 34.38 0 32 20 Zhan et al. [31] 6.11 0.04 71.45 0.24 s 10 5 Chromatic Aberration Ours 24.10 0.61 36.51 0.64 s 10 4
We aim to reconstruct a hyperspectral image $\mathbf{X}\in\mathbb{R}^{H\cdot W\times C}$ of the target scene using a focal stack $\mathbf{Y}\in\mathbb{R}^{H\cdot W\times N}$, where $H$, $W$, and $C$ represent the height, width and channel of the image, respectively. Since $N$ grayscale images in the focal stack are captured by translating the moving lens to different positions $z_{i}$ (more details in Sec. 3.1), we can express the forward model between $\mathbf{X}$ and $\mathbf{Y}$ as follows:
@@
Each block $\mathbf{H}_{i,j}\in\mathbb{R}^{(H+K-1)\cdot(W+K-1)\times H\cdot W}$ is a linear 2D convolutional matrix generated by a PSF $\mathbf{K}(z_{i},\lambda_{j})\in\mathbb{R}^{K\times K}$, associated with the $j$th channel of $\mathbf{X}$ and wavelength $\lambda_{j}$. The binary matrix $\mathbf{C}\in\mathbb{R}^{N\cdot H\cdot W\times N\cdot(H+K-1)\cdot(W+K-1)}$ is used to crop the targeted region of the convolved result (i.e., $\mathbf{H}\mathbf{x}$), ensuring the same pixel count between $\mathbf{x}$ and $\mathbf{y}$. Here, we vectorize the images $\mathbf{X}$ and $\mathbf{Y}$ using column-stacking, defining $\mathbf{x}=\operatorname{vec}(\mathbf{X}):=[\mathbf{X}_{1}^{T},\dots,\mathbf{X}_{C}^{T}]^{T}$ and $\mathbf{y}=\operatorname{vec}(\mathbf{Y})$, where $\mathbf{X}_{j}$ denotes the $j$th column of the matrix $\mathbf{X}$. Here, $\mathbf{X}_{j}$ can also be interpreted as the $j$th channel of the hyperspectral image.
@@ 3.1 System Design
As shown in Fig. 2a, the imaging system comprises an objective lens, a moving lens, and a grayscale sensor. The objective lens and the camera remain fixed while the moving lens is translated to $N$ positions along the optical axis, $z_{1},\cdots,z_{N}$, when capturing the $N$ images in a chromatic focal stack. Given a pre-determined depth of field for the optical system and a set of discrete wavelengths $\lambda_{1},\cdots,\lambda_{N}$, we calibrate the $i$th position of the moving lens $z_{i}$ so that the $i$th wavelength $\lambda_{i}$ is in focus, as shown by the point spread functions in Fig. 2b. Assuming the system measures a target in the depth of field, the captured image $I_{i}$ contains sharp textures corresponding to wavelength $\lambda_{i}$. In contrast, textures of other wavelengths gradually blur out as the wavelengths deviate from $\lambda_{i}$. The total travel distance of the moving lens, $z_{N}-z_{1}$, is very small compared to the effective focal length of the system. Thus, we assume the magnification of the captured images $I_{1},\cdots,I_{N}$ remains constant when capturing the chromatic focal stack.

### v2

@@ Abstract
Hyperspectral cameras face harsh trade-offs between spatial, spectral, and temporal resolution in inherently low-photon conditions. Computational imaging systems break through these trade-offs with compressive sensing, but have required complex optics and/or extensive compute. We present Spectrum from Defocus (SfD), a chromatic focal sweep method that achieves state-of-the-art hyperspectral imaging with only two off-the-shelf lenses, a grayscale sensor, and less than one second of reconstruction time. By capturing a chromatically-aberrated focal stack that preserves nearly all incident light, and reconstructing it with a fast physics-based iterative algorithm, SfD delivers sharp, accurate hyperspectral images. The combination of photon efficiency, optical simplicity, and physical interpretability makes SfD a promising solution for fast, compact, interpretable hyperspectral imaging.
@@ 1 Introduction
Figure 1: The Spectrum from Defocus (SfD) method. (a) Our hardware prototype uses a moving lens to sweep focus through chromatic aberration. (b) SfD achieves state-of-the-art hyperspectral imaging with simple optics and low computational cost. See Table 1 for details. (c) The system captures 5 defocused grayscale images and reconstructs hyperspectral images in under a second.
Hyperspectral imaging (HSI) extends traditional color imaging by capturing detailed spectral signatures across numerous spectral bands, rather than just the three broad RGB bands. This richer spectral information is essential for applications that rely on subtle material differences, such as remote sensing for environmental monitoring [5], medical diagnostics [35], food quality control [15], and industrial inspection [42]. Despite its importance, capturing high-quality spectral images with compact, affordable cameras remains challenging. Hyperspectral sensors suffer from inherently low signal-to-noise ratios because incoming light is divided across many spectral channels, leaving each channel with reduced photon counts. This challenge is amplified in systems that rely on spectral filters, dispersive elements, or coded apertures, all of which block and/or spread light and further reduce photon efficiency. As a result, traditional hyperspectral cameras are particularly fragile in low-light conditions or dynamic scenes that require short exposure times, limiting their practicality.
To mitigate this fragility, compressive spectral imaging systems encode spectral information into spatial measurements, enabling snapshot capture or reduced acquisition times [2]. These systems typically rely on coded optics followed by computational reconstruction. While effective, they often demand heavy computation, involving large matrix operations or iterative solvers. Faster alternatives, such as inverse filtering, degrade reconstruction quality and limit their practical use in real-world scenes [56].
To address the computational challenges of classical reconstruction methods, learning-based approaches have been increasingly explored. These methods often incorporate spectral priors into parameterized architectures to improve efficiency and accuracy. However, these data-driven models risk hallucinating spectral content [17], making them less reliable in tasks where spectral accuracy is critical, such as material analysis or food quality inspection.
To address these limitations in light inefficiency, high computational complexity, and hallucination, we propose a chromatic focal sweep camera consisting of two lenses and a grayscale sensor. By exploiting chromatic aberration as a natural cue for separating wavelengths, our optical system produces a structured sensing matrix that enables computationally efficient and robust hyperspectral reconstruction. Furthermore, this physics-based approach, though boosted by a deep denoising network, mitigates the hallucination issues commonly observed in purely-data-driven methods.
@@
  • Fast computation with an interpretable model that leverages both physics- and data-driven components.
  • A simple optical design consisting of few, off-the-shelf components, which outperforms systems with up to 20 optical elements.
@@ 2 Related Works
Scanning Methods acquire hyperspectral data cubes by sequentially capturing individual 2D slices of the scene. The most straightforward implementation uses band-pass filters or liquid crystal tunable filters to capture one narrow spectral band per frame, gradually scanning across the spectrum [47, 49, 55]. Pushbroom systems take a slightly different approach, capturing a single spatial row at a time through a narrow slit, while dispersive optics spread each row into a full spectrum across the sensor [43]. While these techniques can deliver high spatial and spectral resolution, they suffer from low photon efficiency—since only a fraction of the incoming light is used for each measurement—and long acquisition times, which restricts them to static, well-lit scenes. In contrast, our method passively encodes spectral information into each frame and avoids light-blocking (low light transmission) elements entirely, enabling faster capture with higher photon throughput.
@@
Compressive hyperspectral imaging systems capture optically coded measurements that must be computationally reconstructed to recover the final hyperspectral data cube. Among these, Coded Aperture Snapshot Spectral Imaging (CASSI) [48] was developed to overcome the limitations of sequential scanning systems. A coded aperture is placed either before or after a dispersive prism, encoding spectral signatures into the spatial domain. CASSI systems are typically divided into two categories based on their encoding strategy: spatially-encoded CASSI, which employs a single disperser (SD-CASSI) [48], and spatial-spectral CASSI, which codes information in both spatial and spectral domains (DD-CASSI [18], SS-CASSI [32]). Regardless of the encoding scheme, in all CASSI systems, reconstruction algorithms play a critical role in final image quality. Early approaches relied on optimization-based solvers with hand-crafted priors [48, 4, 33, 53], while more recent approaches introduced learning-based reconstructions [7, 37, 6, 39, 11, 25]. Hybrid methods further combined the system’s sensing matrix with deep learning, either to directly parameterize the solution space [12, 37, 38] or to learn data-driven priors for regularization [50, 52, 57]. We compare our method to two CASSI systems: Choi et al. [12] which uses a hybrid physics- and data-driven reconstruction, and the transformer-based MST [7] which uses spectral-wise self-attention to capture correlations across wavelengths. Despite their success, traditional CASSI systems are often relatively large and mechanically sensitive, making them impractical for compact imaging platforms. In comparison, our system uses only two off-the-shelf lenses and achieves spectral imaging through a focal sweep, significantly reducing hardware complexity.
Diffractive optical elements (DOEs) and metasurfaces
@@
have been designed in tandem with computational reconstruction for compressive spectral sensing [19, 22, 46, 24, 34, 28]. Early work, such as [19], used simple diffraction gratings to simultaneously resolve spatial and spectral information, but these systems suffered from low resolution and bulky optical assemblies. More recently, efforts have focused on using a single DOE, reducing hardware complexity but introducing stronger optical artifacts which are typically handled through end-to-end deep learning reconstructions trained on large spectral datasets [3, 27, 31]. We compare our work with a recent approach, 2in1 Cameras [46], which leverages dual pixels to simultaneously image through a DOE and a conventional lens. Similar to these approaches, our system also relies on spectrally encoded PSFs; however, we achieve this encoding using the naturally occurring chromatic aberration of standard refractive lenses, avoiding the need for custom-fabricated optics.
Figure 2: Optical design. (a) The system consists of a lens pair in which the second lens is translated to five discrete positions, $z_{1},z_{2},\ldots,z_{5}$. At each position $z_{i}$, the captured image $I_{i}$ is focused on a corresponding wavelength $\lambda_{i}$, while other wavelengths appear blurred. Together, these measurements form the chromatic focal stack. (b) Measured point spread functions (PSFs) from the real-world prototype at different wavelengths $\lambda_{i}$ and lens positions $z_{i}$, showing how the focal plane shifts. While calibration PSFs are collected with the help of a narrow-band tunable filter, at deployment time our grayscale sensor measures only the scene-weighted sum of incoming light across all wavelengths at each of the five lens positions.
Other Recent Spectral Computational Imagers highlight the trade-off between optical simplicity and computational cost in spectral imaging. Spectral DiffuserCam (SDC) [41] eliminates conventional lenses by using a diffuser and a specialized sensor equipped with a spectral filter array, analogous to a Bayer pattern. By spreading fine spatial details across multiple pixels, SDC encodes high-frequency information that is later recovered via inverse reconstruction with FISTA. A more recent system, Spectral DefocusCam [16], replaces the diffuser with an achromatic lens, enabling faster acquisition of multiple, less blurry measurements, leading to significantly improved reconstruction quality. Our work shares similarities with Spectral DefocusCam in that both systems encode information via defocus. However, while DefocusCam uses binned super-pixels and exploits blur for spatial super-resolution, our system employs a grayscale sensor and encodes spectral content through spatially varying blur induced by chromatic aberration. Another direction is represented by KRISM [44], which uses a complex optical design to directly measure spectral singular values, thereby shifting most of the computational burden to optical hardware. Our approach strikes a balance between these extremes by combining a simple optical system—using only off-the-shelf parts—with a lightweight, physics-based reconstruction algorithm.
Focal sweeping and stacking form the basis of our compressive sensing approach, in which wavelength-dependent defocus encodes spectral information across a series of focal planes. Variants of this strategy have been explored to extend depth of field [13], enable depth sensing through color-coded apertures [9], reduce chromatic fringing in simple lenses [23], and light field imaging [26]. Spectral reconstruction from focal stacks has been explored in [21], which recovers spatially sparse spectra from PSF center attenuation, and by Zhan et al. [54] to reproduce full hyperspectral images with a naive algorithm. Our robust algorithm produces full hyperspectral images with high fidelity even in the presence of significant noise.
@@ 3 Method
Table 1: Comparison of SOTA hyperspectral imaging systems in terms of reconstruction performance, and computational and optical requirements. Reconstruction quality is benchmarked on 30 images from the Harvard dataset [8] under a brightly lit condition with 5-second total exposure time and optical component effects (see supplement). Timings are reported for an NVIDIA RTX A6000. The count of optical components includes: lenses, apertures, prisms, actuators, SLMs, and sensors, but not control electronics.
Modality Reconstruction Performance Computation and Optics Category PSNR (dB) $\uparrow$ SSIM $\uparrow$ SAM () $\downarrow$ Compute Time FoV () Opt. Components Tunable Filter 4.78 0.08 70.10 0 N.A. 3 Spectral Filter S.DiffuserCam [41] 15.59 0.46 34.98 >> 15 mins 14 3 Spectral Filter S.DefocusCam [16] 21.17 0.61 25.11 39 s 8 4 Spectral Filter MST [7] 30.62 0.92 9.33 1.19 s 10 9 Dispersive Element Choi et al. [12] 23.97 0.68 14.85 >> 15 mins 10 9 Dispersive Element KRISM [44] 29.40 0.91 6.91 0 32 20 Dispersive Element 2in1 Cameras [46] 29.11 0.87 8.72 1.74 s 9 4 DOE Zhan et al. [54] 14.11 0.07 23.95 0.24 s 10 5 Chromatic Aberration Ours 30.81 0.92 7.35 0.64 s 10 4 Chromatic Aberration
We aim to reconstruct a hyperspectral image $\mathbf{X}\in\mathbb{R}^{H\cdot W\times C}$ of the target scene using a focal stack $\mathbf{Y}\in\mathbb{R}^{H\cdot W\times N}$, where $H$, $W$, and $C$ represent the height, width and channel of the image, respectively. Since $N\ll C$ grayscale images in the focal stack are captured by translating the moving lens to different positions $z_{i}$ (more details in Sec. 3.1), we can express the forward model between $\mathbf{X}$ and $\mathbf{Y}$ as follows:
@@
Each block $\mathbf{H}_{i,j}\in\mathbb{R}^{(H+K-1)\cdot(W+K-1)\times H\cdot W}$ is a linear 2D convolutional matrix generated by a PSF $\mathbf{K}(z_{i},\lambda_{j})\in\mathbb{R}^{K\times K}$, associated with the $j$th channel of $\mathbf{X}$ and wavelength $\lambda_{j}$. The binary matrix $\mathbf{C}\in\mathbb{R}^{N\cdot H\cdot W\times N\cdot(H+K-1)\cdot(W+K-1)}$ is used to crop the targeted region of the convolved result (i.e., $\mathbf{H}\mathbf{x}$), ensuring the same pixel count between $\mathbf{x}$ and $\mathbf{y}$. We vectorize the images $\mathbf{X}$ and $\mathbf{Y}$, defining $\mathbf{x}=\operatorname{vec}(\mathbf{X}):=[\mathbf{X}_{1}^{T},\dots,\mathbf{X}_{C}^{T}]^{T}$ and $\mathbf{y}=\operatorname{vec}(\mathbf{Y})$, where $\mathbf{X}_{j}$ denotes the $j$th column of the matrix $\mathbf{X}$. Here, $\mathbf{X}_{j}$ can also be interpreted as the $j$th channel of the hyperspectral image. Unless otherwise specified $N=5$ is assumed throughout (see supplement).
@@ 3.1 System Design
As shown in Fig. 2a, the imaging system consists of an objective lens, a movable lens, and a grayscale sensor. The objective lens and the camera remain fixed while the second lens is translated to $N$ positions along the optical axis, $z_{1},\cdots,z_{N}$, to capture the $N$ images, forming a chromatic focal stack. Each lens position $z_{i}$ is calibrated so that the corresponding wavelength $\lambda_{i}$ is in focus, as illustrated by the measured point spread functions (PSFs) in Fig. 2b. When the scene lies within the system’s depth of field, the captured image $I_{i}$ contains sharp structures at wavelength $\lambda_{i}$, while other wavelengths appear increasingly blurred as they deviate from $\lambda_{i}$. The total travel distance of the moving lens, $z_{N}-z_{1}$, is small compared to the system’s effective focal length, so we assume the magnification of $I_{1},\ldots,I_{N}$ remains constant across the stack.